# Native Language Identification in Turkish: L1 Influence of Arabic, Persian, and Albanian

### v5

@@ 1 Introduction
Native Language Identification (NLI) is the task of automatically identifying the native language (L1) of an individual based on their linguistic productions in another language (L2). The underlying hypothesis is that the L1 influences learners’ second language writing as a result of the language transfer effect (Yu and Odlin, 2016). It is used for a variety of purposes including forensics applications in cybercrime (Perkins, 2021) and second language acquisition (Swanson and Charniak, 2014).
Research in NLI is mainly conducted with learner corpora, which comprise collections of writings by individuals learning a new language. These writings are annotated with metadata such as the author’s native language (L1) or their fluency level. Recent NLI studies on languages other than English include Portuguese (Malmasi et al., 2018), Arabic (Malmasi and Dras, 2014a), and Chinese (Malmasi and Dras, 2014b). The learner corpus is the backbone of NLI research, which means that extending research to a novel language depends on acquiring the appropriate learner corpora for that language. In the past, studies have focused on L2 English because of the prominence of this language in language research and the relatively large amount of data available. To the best of our knowledge, this study presents the first comprehensive NLI experiments on L2 Turkish. We employ the recently constructed Turkish Learner Corpus (TLC) (Golynskaia, 2022) and investigate widely used syntactic features for NLI. The remainder of this paper is organised as follows: Section 2 reviews related work, Section 3 and 4 describe our data and methodology, Section 5 presents the results and our discussion, and Section 6 concludes the paper.
@@ 2 Related Work
NLI is typically modeled as a supervised multi-class classification task. In this experimental design, the individual writings of learners are used to train a model while the author’s L1 information serves as class labels. A variety of feature types at the syntactic and lexical levels have been studied to capture distinct characteristics of the language interference phenomenon: spelling errors, word and lemma n-grams, dependency parsing, and morphosyntax. A more detailed review of linguistic features can be found in two shared task reports on the NLI task organised in 2013 and 2017 (Tetreault et al., 2013; Malmasi et al., 2017).
In recent years, there has been increased experimentation with deep learning methods, including pre-trained encoder models such as BERT (Steinbakken and Gambäck, 2020) and decoder models such as GPT-2 (Lotfi et al., 2020). While these models outperformed the state-of-the-art performance achieved by feature-based stacked classifiers (Malmasi and Dras, 2018), questions about their interpretability, inherent biases, and practical shortcomings in industrial applications remain unexplored. Traditional methods based on hand-crafted features continue to be preferred in many implementations due to their simplicity in training and resource efficiency. Within this context, Uluslu and Schneider (2022) approached the NLI scalability problem in the context of cybercrime through model compression and inference speed.
@@ 3 Data
In this study, we use data from the TLC (Golynskaia, 2022). TLC is a learner corpus composed of the writings of learners of Turkish. These texts are essays written as part of a test of Turkish as a secondary language. While all learners in this study had completed a B1 level Turkish course, we observed a lack of standardized CEFR-based proficiency assessment in the material. Each text includes additional metadata such as the nationality of the author and the genre of the text. The corpus also includes error codes and corrections, although we do not make use of this information. In the second version of the NLI corpus, we removed learner texts from Azerbaijan (due to its close similarity to Turkish) and Afghanistan (due to the difficulty in determining a specific L1 among Pashto, Dari, and Persian).
@@ 4.1 Classifier
For this NLI task, we employ a standard supervised multi-class classification approach. Following previous research (Gebre et al., 2013), we use a linear Support Vector Machine (SVM) classifier. The input features are n-grams weighted with the Term Frequency-Inverse Document Frequency (TF-IDF) scheme, which our preliminary experiments showed outperformed simple relative frequencies. To optimize the model, we performed a grid search for the regularization parameter C over the range $10^{-3}$, $10^{-1}$] and set max_iter=2000 to ensure convergence. The model’s performance plateaued at the upper end of this range, so we selected $C=1.0$ for our final experiments.
@@ 4.3 Linguistic Features
We focus only on content-independent features, in particular syntactic features, following the example of studies on NLI in other languages (Malmasi et al., 2015). Due to the imbalance in the topic distribution in the TLC corpus, we decided not to include lexical features such as word n-grams and contextualized word embeddings in our study. For example, only students with Persian L1 were asked to respond to prompts specifically about happiness and time. This can result in the classifier associating these topics with the languages, rather than discerning the characteristics inherent to the language, thereby introducing a confounding variable to the task (Brooke and Hirst, 2013). By focusing on syntactic features, we aim to capture the underlying syntactic influence of the L1 on its L2 writing independently of the content (Malmasi et al., 2018). We experimented using a combination of two syntactic features: part-of-speech n-grams and function words.
@@
Function words are content-independent words, including prepositions, articles, and auxiliary verbs, that play a crucial role in conveying grammatical relationships between words. It is often challenging for L2 speakers to use the appropriate function words and production errors may be due to the influence of their L1 (Schneider and Gilquin, 2016). These function words are recognised as valuable features for the NLI task. We extracted 120 Turkish words from different grammatical categories including delexicalised verbs. However, it’s worth noting that many grammatical aspects, which are morphologically expressed in Turkish, may not be as strongly captured as they would be in English.
Part-of-Speech (POS) tags are linguistic categories or word classes that signify the syntactic role of each word in a sentence. They include basic categories such as verbs, nouns, and adjectives. Assigning POS tags to words in a text introduces a level of linguistic abstraction, meaning that we can work with the underlying structure rather than the content. We use the Turkish NLP toolkit222turkish-nlp-suite/tr_core_news_trf from SpaCy (Honnibal et al., 2020) to extract universal POS tags, from which we create n-grams of sizes 1 to 3. These n-grams serve to capture preferences for specific word classes and their localized ordering patterns. Our experiments indicated that sequences of order 4 or higher lead to lower accuracy due to the limited size of our corpus. Therefore, we excluded such higher-order n-grams from our analysis. When both features are combined, function words retain their form instead of being replaced by their POS tags. Figure 1 provides an example of these extraction steps.
@@ 7 Conclusion
Our analysis brings attention to two potential limitations. Firstly, the size of our corpus, while sufficient for achieving results significantly above the baseline, is relatively limited compared to large-scale NLI studies for Portuguese and Norwegian (Malmasi et al., 2018, 2015). It is, however, comparable in scale to the corpus used for the first Arabic NLI study (Malmasi and Dras, 2014a). Secondly, we acknowledge that our parser might not be entirely suitable for learner language, which could introduce additional noise into the feature space Van Rooy and Schäfer (2002). However, we chose to use the spaCy transformer-based model, as a recent comparative study has shown it to be relatively robust for processing learner language compared to alternatives (Kyle and Eguchi, 2024), which helps mitigate this concern.

### v6

# Turkish Native Language Identification V2: L1 Influence of Arabic, Persian, and Albanian

@@ 1 Introduction
Native Language Identification (NLI) is the task of automatically identifying the native language (L1) of an individual based on their linguistic productions in another language (L2). The underlying hypothesis is that the L1 influences learners’ second language writing as a result of the language transfer effect (yu2016new). It is used for a variety of purposes including forensics applications in cybercrime (perkins2021application) and second language acquisition (swanson2014data).
Research in NLI is mainly conducted with learner corpora, which comprise collections of writings by individuals learning a new language. These writings are annotated with metadata such as the author’s native language (L1) or their fluency level. Recent NLI studies on languages other than English include Portuguese (malmasi2018portuguese), Arabic (malmasi2014arabic), and Chinese (malmasi2014chinese). The learner corpus is the backbone of NLI research, which means that extending research to a novel language depends on acquiring the appropriate learner corpora for that language. In the past, studies have focused on L2 English because of the prominence of this language in language research and the relatively large amount of data available. To the best of our knowledge, this study presents the first comprehensive NLI experiments on L2 Turkish. We employ the recently constructed Turkish Learner Corpus (TLC) (anna2022error) and investigate widely used syntactic features for NLI. The remainder of this paper is organised as follows: Section 2 reviews related work, Section 3 and 4 describe our data and methodology, Section 5 presents the results and our discussion, and Section 6 concludes the paper.
@@ 2 Related Work
NLI is typically modeled as a supervised multi-class classification task. In this experimental design, the individual writings of learners are used to train a model while the author’s L1 information serves as class labels. A variety of feature types at the syntactic and lexical levels have been studied to capture distinct characteristics of the language interference phenomenon: spelling errors, word and lemma n-grams, dependency parsing, and morphosyntax. A more detailed review of linguistic features can be found in two shared task reports on the NLI task organised in 2013 and 2017 (tetreault2013report; malmasi2017report).
In recent years, there has been increased experimentation with deep learning methods, including pre-trained encoder models such as BERT (steinbakken2020native) and decoder models such as GPT-2 (lotfi2020deep). While these models outperformed the state-of-the-art performance achieved by feature-based stacked classifiers (malmasi2018native), questions about their interpretability, inherent biases, and practical shortcomings in industrial applications remain unexplored. Traditional methods based on hand-crafted features continue to be preferred in many implementations due to their simplicity in training and resource efficiency. Within this context, uluslu2022scaling approached the NLI scalability problem in the context of cybercrime through model compression and inference speed.
@@ 3 Data
In this study, we use data from the TLC (anna2022error). TLC is a learner corpus composed of the writings of learners of Turkish. These texts are essays written as part of a test of Turkish as a secondary language. While all learners in this study had completed a B1 level Turkish course, we observed a lack of standardized CEFR-based proficiency assessment in the material. Each text includes additional metadata such as the nationality of the author and the genre of the text. The corpus also includes error codes and corrections, although we do not make use of this information. In the second version of the NLI corpus, we removed learner texts from Azerbaijan (due to its close similarity to Turkish) and Afghanistan (due to the difficulty in determining a specific L1 among Pashto, Dari, and Persian).
@@ 4.1 Classifier
For this NLI task, we employ a standard supervised multi-class classification approach. Following previous research (gebre2013improving), we use a linear Support Vector Machine (SVM) classifier. The input features are n-grams weighted with the Term Frequency-Inverse Document Frequency (TF-IDF) scheme, which our preliminary experiments showed outperformed simple relative frequencies. To optimize the model, we performed a grid search for the regularization parameter C over the range $10^{-3}$, $10^{-1}$] and set max_iter=2000 to ensure convergence. The model’s performance plateaued at the upper end of this range, so we selected $C=1.0$ for our final experiments.
@@ 4.3 Linguistic Features
We focus only on content-independent features, in particular syntactic features, following the example of studies on NLI in other languages (malmasi2015norwegian). Due to the imbalance in the topic distribution in the TLC corpus, we decided not to include lexical features such as word n-grams and contextualized word embeddings in our study. For example, only students with Persian L1 were asked to respond to prompts specifically about happiness and time. This can result in the classifier associating these topics with the languages, rather than discerning the characteristics inherent to the language, thereby introducing a confounding variable to the task (brooke2013native). By focusing on syntactic features, we aim to capture the underlying syntactic influence of the L1 on its L2 writing independently of the content (malmasi2018portuguese). We experimented using a combination of two syntactic features: part-of-speech n-grams and function words.
@@
Function words are content-independent words, including prepositions, articles, and auxiliary verbs, that play a crucial role in conveying grammatical relationships between words. It is often challenging for L2 speakers to use the appropriate function words and production errors may be due to the influence of their L1 (schneider2016detecting). These function words are recognised as valuable features for the NLI task. We extracted 120 Turkish words from different grammatical categories including delexicalised verbs. However, it’s worth noting that many grammatical aspects, which are morphologically expressed in Turkish, may not be as strongly captured as they would be in English.
Part-of-Speech (POS) tags are linguistic categories or word classes that signify the syntactic role of each word in a sentence. They include basic categories such as verbs, nouns, and adjectives. Assigning POS tags to words in a text introduces a level of linguistic abstraction, meaning that we can work with the underlying structure rather than the content. We use the Turkish NLP toolkit222turkish-nlp-suite/tr_core_news_trf from SpaCy (honnibal2020spacy) to extract universal POS tags, from which we create n-grams of sizes 1 to 3. These n-grams serve to capture preferences for specific word classes and their localized ordering patterns. Our experiments indicated that sequences of order 4 or higher lead to lower accuracy due to the limited size of our corpus. Therefore, we excluded such higher-order n-grams from our analysis. When both features are combined, function words retain their form instead of being replaced by their POS tags. Figure 1 provides an example of these extraction steps.
@@ 7 Conclusion
Our analysis brings attention to two potential limitations. Firstly, the size of our corpus, while sufficient for achieving results significantly above the baseline, is relatively limited compared to large-scale NLI studies for Portuguese and Norwegian (malmasi2018portuguese; malmasi2015norwegian). It is, however, comparable in scale to the corpus used for the first Arabic NLI study (malmasi2014arabic). Secondly, we acknowledge that our parser might not be entirely suitable for learner language, which could introduce additional noise into the feature space van2002effect. However, we chose to use the spaCy transformer-based model, as a recent comparative study has shown it to be relatively robust for processing learner language compared to alternatives (kyle2024evaluating), which helps mitigate this concern.